OpenStudy (anonymous):

An oil company distributes oil in a metal can shaped like a cylinder that has an actual radius of 5.1cm and a height of 15.1cm. A worker incorrectly measured the radius as 5cm and the height as 15cm. Determine the relative error in calculating the surface area, to the nearest thousandth.

OpenStudy (amistre64):

compare volumes is what im thinking

OpenStudy (amistre64):

or surface areas as this case may be

OpenStudy (amistre64):

i got no idea what a "relative" error is tho; terminology still eludes me from time to time

OpenStudy (anonymous):

Thanks! I'll try to figure it out

OpenStudy (amistre64):

2pi r h + 2pi r^2 is the formula for surface area, that i think i recall correctly

OpenStudy (amistre64):

2pi r(h+r)

OpenStudy (anonymous):

SA = 2π r h + 2π r² = 2π r ( h + r ) = 2π *5 ( 10 + 5) = 471.24 cm² SA' = 2π (r' h + r h') + 4π r r' = 2π ( 10 *.1 + 5 * .1) + 4π * 5 * .1 = 2π *1.5 + 2π = 15.7 cm² => Relative error: 15.7/ 471.24 = .0333 = 3%